# CREST-Search: Comprehensive Red-teaming for Evaluating Safety Threats in Large Language Models Powered by Web Search

### v1

@@
Abstract.
Large Language Models (LLMs) have rapidly advanced the state of natural language processing, powering applications such as dialogue systems, summarization, and question answering. Despite their remarkable capabilities, LLMs struggle to adapt to specialized domains and capture emerging facts or evolving contexts beyond the cutoff of their training data. Driven by these limitations, web search has been integrated into LLM systems, enabling the models to access and synthesize information directly from the open Internet. Such connection to the open web, however, magnifies long-standing safety concerns. LLMs are already vulnerable to adversarial prompts; when paired with untrusted online content, these risks are amplified, exposing users to more severe and diverse safety breaches. Such challenges cast serious doubt on the reliability and safety of deploying LLM applications with web search in practice.
This paper investigates red teaming for LLMs with web search, aiming to systematically uncover vulnerabilities in such services. Existing red teaming solutions designed for standalone LLMs are insufficient in this context due to the intricate search workflow, which integrates reasoning, generation, and external tool use. To address these challenges, we introduce CREST-Search, a novel red-teaming framework tailored for LLMs with web search. Guided by three pioneering attack strategies, CREST-Search leverages in-context learning to generate initial test search queries and iteratively refines them with feedback to enhance their adversarial strength, thereby exposing the risk during the search process. Another contribution of CREST-Search is the construct of a search-specific harmful dataset (WebSearch-Harm), It can be used to fine-tune a base LLM into a specialized red-teaming model, which further enhances the effectiveness of iterative refinement while reducing cost and consumption time. Our experiments demonstrate that CREST-Search can effectively bypass safety filters and systematically expose vulnerabilities in modern web search-based LLM systems. These underscore the necessity of specialized solutions to support the development of trustworthy and reliable search models.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Large Language Models (LLMs), such as OpenAI GPT-4o (OpenAI, 2024a), Gemini (Team, 2025), and LLaMA (Dubey et al., 2024), have demonstrated remarkable capabilities across a wide range of natural language processing (NLP) tasks, including reasoning (Wei et al., 2022), question answering (Kwiatkowski et al., 2019), and code generation (Chen et al., 2021). Beyond NLP, their impact has rapidly expanded to high-stakes domains such as education (Wang et al., 2023), healthcare (Thirunavukarasu et al., 2023), and finance (Yang et al., 2024), where they increasingly serve as critical components in decision support systems. Despite these advances, LLMs suffer from an inherent limitation: their knowledge is bounded by the training cutoff date, making them unable to incorporate newly emerging facts, evolving events, or rapidly changing domain-specific knowledge (Ji et al., 2023). To address this, web search functionality has been integrated into LLMs (OpenAI, 2024a; DeepMind, 2024), enabling models to dynamically access external knowledge from the open Internet. Compared with traditional keyword-based search engines, LLMs with web search allow users to query in natural language and generate concise answers by automatically aggregating information across multiple sources, greatly improving accessibility and usability (Nakano et al., 2021). A survey conducted by Rowlands et.al (Rowlands, 2025) reveals an increasing trend that more and more people prefer to use generative AI tools like ChatGPT to replace the traditional search engines for information retrieval.
Despite their impressive capabilities, LLMs continue to suffer from fundamental safety vulnerabilities. A growing body of research has demonstrated that adversaries can craft carefully engineered queries to circumvent safety filters and alignment mechanisms, ultimately coercing models into producing harmful or undesirable outputs. Prominent attack vectors include jailbreak attacks (Yi et al., 2024), prompt injection (Kumar et al., 2023), and adversarial manipulation (Yao et al., 2023). Beyond academic findings, policymakers have also raised alarms: the UK government’s official discussion paper (UK, 2023) cautions that generative AI systems may significantly accelerate the production and dissemination of harmful content, with the potential to erode public trust, distort democratic processes, and escalate both societal and national security risks. Similarly, the European Union’s AI Act (Smuha, 2025) highlights the need for stringent oversight and safeguards for high-risk AI applications, underscoring the broader policy recognition of safety risks associated with large-scale generative models.
@@ 1. Introduction
The integration of web search into LLMs further introduces new safety risks. One the one hand, different from standalone offline LLMs whose responses are confined to their training data, online LLM-based systems invoke search engines to generate answers. They may retrieve information from uncontrolled and heterogeneous webpages in the wild, which could contain malicious or low-credibility content that cannot be reliably filtered by built-in safety mechanisms. As a result, these systems carry an elevated risk of exposing users to misinformation, hate speech, and other harmful material. On the other hand, prior works have demonstrated that traditional search engines are susceptible to adversarial attacks. Malicious users can subvert the indexing and retrieval mechanisms of search engines such as Google and Bing by inserting imperceptible text-encoding perturbations into benign search queries (Boucher et al., 2023). Complementary measurement studies (Joslin et al., 2019) further reveal that a typo coinciding with a legitimate word in the same or another language can bypass auto-correction and enable blacklisted or illicit websites to achieve first-page ranking. Therefore, LLMs with web search can inherit or even amplify such risks. For instance, Luo et.al (Luo et al., 2025) demonstrated that adversaries can exploit search-engine-optimization techniques to manipulate retrieval results, causing AI-driven systems to reference pre-published malicious webpages as authoritative sources or synthesize and output harmful code.
Therefore, it is important to systematically and comprehensively understand and examine the safety vulnerabilities of existing LLM systems augmented by web search, as such assessments offer valuable feedback to developers for strengthening safety mechanisms and improving system robustness. Red-teaming, originating from traditional cybersecurity scenarios, is a promising strategy to achieve the above purpose. Its basic idea is to simulate real-world attacks on the target system and identify its vulnerabilities that could be exploited to achieve actual damages. For real-world applicability, red-teaming must be automated, efficient, and diversity-aware: the system should rapidly synthesize a wide variety of adversarial cases that cover different attack surfaces, enabling scalable, repeatable testing and actionable insights for developers. With the rise of LLMs, researchers have proposed new red-teaming methods to assess the potential of these models in generating inappropriate content such as harmful, bias or privacy output (Perez et al., 2022; Ganguli et al., 2022; An et al., 2025). For instance, testers can craft a series of jailbreak prompts (Ge et al., 2023) to make the LLMs generate toxic or disallowed content or steal the private personal data (Kim et al., 2023) from the models by multi-turn inference.
Existing red-teaming methods exclusively focus on LLMs that are standalone or equipped with Retrieval-Augmented Generation (RAG). They exhibit significant limitations when applied to the safety assessment of LLMs with web search. (1) Limited coverage for risk discovery. When the LLM exploits the web search to produce the responses, safety vulnerabilities may emerge at multiple stages of the generation pipeline, including information retrieval, filtering, answer generation, and citation. Conventional red-teaming primarily focuses on the safety of the generated answers, but overlooks critical citation risks existing in LLMs with web search because vanilla LLMs do not incorporate the stage of retrieving online information and citing external webpages. This primary difference makes conventional red-teaming ineffective to directly transfer the methods developed for standalone LLMs to the setting involving web search. (2) Complex and opaque generation pipeline. Real-world LLM systems integrated with web search are typically black-box online services provided by commercial companies. Their internals of retrieval, generation, citation, and safety filtering remain opaque, leading to stochastic and unpredictable behaviors. Users can only use this service by invoking the API endpoint. However, existing red-teaming approaches mainly research on the open-source models. Therefore, researchers can conduct the test under white-box or gray-box settings where the pre-filtering or post-filtering is removed. This substantially reduces the difficulty of testing, which in turn diminishes their effectiveness when evaluating on commercial black-box models (Li et al., 2024; Chen et al., [n. d.]). The above gaps call for an efficient and effective methodology dedicated to web-search LLMs, which operate purely from inputs and outputs, with high scalability and reproducibility.
@@
We propose CREST-Search, a novel Comprehensive Red-teaming approach for Evaluating Safety Threats in LLMs with Web Search. CREST-Search adopts a three-stage pipeline, consisting of adversarial search query generation, web search execution and risk evaluation, and adversarial search query refinement. It further incorporates several innovative techniques to address the above challenges. (1) To uncover the unique citation risk in the LLMs with web search, we propose three novel adversarial test search query generation strategies, which are specifically tailored to this testing scenario. We further introduce a risk classification that considers responses, citation, as well as their combinations to systematically report the test results. We then employ a set of metrics to comprehensively evaluate the quality of adversarial queries, including overall risk detection rate, query toxicity, and diversity. These metrics allow us to assess whether CREST-Search can generate adversarial search queries that effectively expose risks while keeping the queries non-trivially offensive and maintaining controlled self-similarity. (2) To effectively test under black-box online constraints, we leverage a dedicated red-teaming model for adversarial query generation. This model is built by supervised fine-tuning on our crafted web-search–specific harmful dataset (WebSearch-Harm), which includes conversations annotated with generation strategies, harmful content categories, and successful attack queries. Then, we incorporate the model with the judgment-feedback refinement loop to optimize weak queries via in-context learning. This design can offer the advantage of reducing the refinement rounds and API costs while improving risk detection under a fixed budget. With these innovations, CREST-Search provides a unique, systematic, and scalable red-teaming framework for LLMs with web search, enabling citation risk identification and efficient adversarial query generation under realistic black-box constraints.
@@
To evaluate CREST-Search, we conduct comprehensive experiments with repeated trials to obtain stable results on 4 commercial LLMs with web search, covering 3 test case generation strategies and 5 categories of harmful content. Compared with baselines, CREST-Search achieves a higher risk detection rate of 80.5% across the 4 LLMs. Through detailed analysis, we find that CREST-Search can successfully expose vulnerabilities across all five categories, underscoring the utility of our proposed test case generation strategies. Our ablation study further highlights the contribution of each component of CREST-Search: both the refinement mechanism and the fine-tuned red-teaming model enhance the optimization efficiency and improve the risk detection rate. These findings demonstrate that CREST-Search serves as a robust red-teaming framework for assessing the robustness, safety, and reliability of LLMs with web search. Our main contributions are as follows:
@@
    We construct a diverse specialized dataset for red-teaming model fine-tuning, named WebSearch-Harm. Compared with adversarial queries generated by the vanilla models, cases generated by the fine-tuned red-teaming models effectively increase the risk detection rate, while reducing the refinement cost and time.
@@
    We propose an innovative red-teaming framework, incorporating three attack strategies, that comprehensively assesses the safety vulnerabilities in the LLMs with web search under the black-box settings, reflecting real-world application scenarios.
@@
    By analyzing the vulnerabilities uncovered through our red-teaming framework, we provide suggestions for enhancing the safety, robustness, and trustworthiness of LLMs with web search.
2. Related Work
2.1. LLMs Enhanced by Web Search
LLMs, represented by GPT-4o (OpenAI, 2024a), Gemini 2.0 Flash (Team, 2025), and LLaMA (Dubey et al., 2024), have achieved remarkable advances in natural language understanding and generation. They demonstrate strong performance across tasks like reasoning (Wei et al., 2022), coding (Chen et al., 2021), and question answering (Kwiatkowski et al., 2019). Beyond NLP, LLMs are increasingly applied in domains such as education (Wang et al., 2023), healthcare (Thirunavukarasu et al., 2023), and finance (Yang et al., 2024), highlighting their growing societal and industrial impact.
LLMs face an inherent limitation: their knowledge is constrained by the fixed cut-off date of their training data. To alleviate this, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has been widely adopted to retrieve relevant information from a maintained external corpus and integrate it into the response generation process. While this is effective in specific domains, RAG typically requires the costly construction (e.g. large-scale data collection, cleaning, and annotation) and continuous maintenance and update, especially in rapidly evolving domains of the specific knowledge base. This poses heavy burdens for personal users or small organizations.
Alternatively, recent works begin to integrate the web search functionality into LLMs, such as GPT-4o-search (OpenAI, 2025) and Gemini-2.5-flash-search (Google, 2025). This allows LLMs to dynamically retrieve information from the open web, thereby generating the answers based on the update-to-date content. Compared with traditional keyword-based search engines like Google (Hillis et al., 2012), LLMs with web search have two advantages. First, they allow users to build search queries with natural language without specialized query syntax or handcrafted search operators. For example, a complex condition search, consisting of multiple manual filters or “search hacks” in a traditional search engine, can be expressed as a single sentence for LLMs. This is friendly and convenient for rookie users not familiar with professional searching skills to precisely access the information they need. Second, when searching the relevant sources, LLMs can produce concise answers by automatically aggregating and summarizing relevant content across multiple webpages. Such capability reduces the effort of browsing and cross-referencing different sources and greatly improves user accessibility.
However, this integration also introduces new security and trustworthiness problems. Since LLMs’ reasoning process partly relies on the content from the Internet, malicious or low-credibility webpages can be retrieved and cited as references, leading to the spread of misinformation or harmful content. There is a pressing need for a systematic investigation into the safety assessment of LLMs augmented with web search.
2.2. Adversarial Attacks Against LLMs
@@
Owing to limitations in training data quality and imperfect safety alignment, LLMs remain susceptible to issues such as hallucination, social and cultural bias, and generation of harmful content. Malicious users can craft carefully-designed prompts or poisoning context, which can effectively bypass the safeguard mechanisms of LLMs, and mislead them to produce inappropriate content.
Numerous attack techniques, such as prompt injection, jailbreak prompting, and adversarial prompt perturbation, have been proposed to exploit these vulnerabilities (Kumar et al., 2023; Yi et al., 2024; Yao et al., 2023). For example, Kumar et.al (Kumar et al., 2023) systematically investigate three adversarial prompting scenarios: suffix attacks, insertion attacks, and infusion attacks. These attacks inject the malicious token into different positions of the user prompt to induce the LLMs to generate harmful content. Yao et.al (Yao et al., 2023) introduce an automated hallucination attack using gradient-based token replacement, which successfully induces fabricated outputs in LLMs through weak semantic and out-of-distribution prompts. These attacks significantly threaten various real-world applications built upon LLMs, including web search.
2.3. Red-teaming LLMs
Building on the above security concerns, red-teaming has emerged as a pivotal strategy for systematically probing the vulnerabilities of LLMs. Red-teaming refers to the deliberate design of adversarial prompts to elicit undesired or harmful outputs, thereby stress-testing a model’s safety alignment and uncovering potential risks (Ganguli et al., 2022). This process can provide insightful guidance for further safety refinement. Many LLM service providers (e.g., OpenAI, Anthropic) (Longpre et al., 2024; Feffer et al., 2024) have adopted red-teaming as an essential step before releasing LLM-based products to the public.
@@ 2.3. Red-teaming LLMs
Researchers have proposed several approaches to red-teaming LLM-based systems (Ganguli et al., 2022; Perez et al., 2022). For instance, Ge et.al (Ge et al., 2023) propose an iterative framework, which leverages an adversarial LLM to generate evolving jailbreak prompts. Those prompts, paired with corrected safe responses, are further used to fine-tune the target LLM, achieving scalable and automated safety enhancement for LLMs. An et. al (An et al., 2025) conduct a systematic analysis towards the safety of RAG for LLMs, demonstrating that even with safe models and safe corpora, RAG can introduce new unsafe behaviors.
@@
Figure 1. Overview of CREST-Search, consisting of three main phases. (1) Adversarial search queries generation: it generates the adversarial search queries based on various harmful content categories and strategies; (2) Web search execution and risk evaluation: it executes the search query and evaluates the toxicity of the cited webpages; (3) Adversarial search queries refinement: it optimizes the query based on the judgment or terminates the round.
3. Preliminary
3.1. Problem Statement

### v2

# When Search Goes Wrong: Red-Teaming Web-Augmented Large Language Models

@@
Abstract
Large Language Models (LLMs) have been augmented with web search to overcome the limitations of the static knowledge boundary by accessing up-to-date information from the open Internet. While this integration enhances model capability, it also introduces a distinct safety threat surface: the retrieval and citation process has the potential risk of exposing users to harmful or low-credibility web content. Existing red-teaming methods are largely designed for standalone LLMs as they primarily focus on unsafe generation, ignoring risks emerging from the complex search workflow. To address this gap, we propose CREST-Search, a pioneering red-teaming framework for LLMs with web search. The cornerstone of CREST-Search is three novel attack strategies that generate seemingly benign search queries yet induce unsafe citations. It also employs an iterative in-context refinement mechanism to strengthen adversarial effectiveness under black-box constraints. In addition, we construct a search-specific harmful dataset, WebSearch-Harm, which enables fine-tuning a specialized red-teaming model to improve query quality. Our experiments demonstrate that CREST-Search can effectively bypass safety filters and systematically expose vulnerabilities in web search-based LLM systems, underscoring the necessity of the development of robust search models.
1 Introduction
Figure 1: Citation risk in LLMs powered by web search.
Large Language Models (LLMs), such as GPT-4o (OpenAI, 2024a), Gemini (Team, 2025), and LLaMA (Dubey and others, 2024), have achieved remarkable performance across a wide range of natural language processing tasks (Wei et al., 2022; Kwiatkowski et al., 2019; Chen et al., 2021) and are increasingly deployed in high-stakes scenarios  (Wang and others, 2023; Yang et al., 2024; Thirunavukarasu and others, 2023)) to make decisions. However, their knowledge is inherently constrained by training cutoff dates, limiting their ability to incorporate newly emerging information (Ji et al., 2023). To address this gap, recent LLM systems have been augmented with web search (OpenAI, 2024a; DeepMind, 2024), enabling models to dynamically retrieve up-to-date information from the open Internet. Compared to traditional search engines, these systems can better understand the semantics of search queries and synthesize concise answers by aggregating information from multiple sources (Nakano et al., 2021).
Despite these advances, LLMs continue to suffer from fundamental safety vulnerabilities. Prior studies show that adversaries can exploit weaknesses in model alignment through adversarial attack techniques (Kumar et al., 2023; Yao et al., 2023) or jailbreak patterns (Yi et al., 2024) to induce harmful behaviors. These risks are not only a technical concern but have also attracted regulatory scrutiny. Policymakers (UK, 2023) warn that generative AI systems may accelerate the dissemination of harmful content and pose broader societal and security threats (Smuha, 2025).
@@ 1 Introduction
The integration of web search into LLMs further introduces new safety risks (Figure 1). On the one hand, unlike standalone LLMs whose outputs are bounded by internal model knowledge, search-enabled LLMs retrieve and aggregate information from heterogeneous and uncontrolled webpages in the wild. Harmful or low-credibility content that bypasses built-in safety mechanisms can be indirectly introduced through external citations. As a result, these systems carry an elevated risk of exposing users to misinformation, hate speech, and other harmful material. On the other hand, LLMs with web search may inherit and potentially amplify vulnerabilities known in traditional search engines. Prior work shows that adversaries can manipulate search indexing and retrieval through techniques such as imperceptible text-encoding perturbations (Boucher et al., 2023) or linguistic collisions (Joslin et al., 2019), allowing malicious or blacklisted webpages to surface in high-ranking positions among search results. Recent studies (Luo et al., 2025; Dong et al., 2025) further demonstrate that such attacks can be exploited in LLM-based search systems, where models cite pre-published malicious webpages deliberately created by adversaries as authoritative sources to synthesize outputs.
Red-teaming serves as a promising approach to evaluate the safety of LLMs under adversarial prompting (Perez et al., 2022; Ganguli et al., 2022; An et al., 2025; Ge et al., 2023) or multi-turn adversarial interactions techniques  (Kim et al., 2023). However, existing methods primarily focus on eliciting unsafe generations from standalone LLMs or those equipped with Retrieval-Augmented Generation (RAG), exhibiting fundamental limitations when applied to LLMs with web search. (1) Limited coverage for risk discovery: safety risks in search-enabled LLMs can emerge at multiple stages of the generation pipeline, including information retrieval, filtering, answer generation, and citation. Conventional red-teaming methods evaluate only the generated responses from the models, overlooking these citation risks that are unique to LLMs with web search. This primary difference makes them ineffective to directly transfer to the setting involving web search. (2) Complex and opaque generation pipeline: real-world search-enabled LLMs are typically deployed as black-box commercial services, with opaque retrieval, filtering, and citation mechanisms. In contrast, most existing red-teaming approaches are developed and evaluated on open-source models under white-box or gray-box settings (Li et al., 2024; Chen et al., ), underestimating the difficulty of probing safety failures in realistic deployment scenarios. These limitations underscore the need for a scalable, black-box red-teaming methodology specifically for such models.
@@
We propose CREST-Search, a novel Comprehensive Red-teaming approach for Evaluating Safety Threats in LLMs with Web Search. It incorporates several innovative techniques to address the above challenges. (1) To uncover the unique citation risk in the LLMs with web search, we propose three novel adversarial search query generation strategies. We further introduce a unified risk classification that considers responses, citation and their combinations to systematically report test results. We then employ a set of metrics to guarantee the quality of adversarial search queries that effectively expose risks while keeping the queries non-trivially offensive and maintaining controlled self-similarity. (2) To effectively test under black-box online constraints, we leverage a dedicated red-teaming model for adversarial query generation. This model is built by supervised fine-tuning on our crafted web-search–specific harmful dataset (WebSearch-Harm), which includes conversations annotated with generation strategies, harmful content categories, and successful attack queries. Then, we incorporate the model with the judgment-feedback refinement loop to optimize weak queries via in-context learning.
@@
To evaluate CREST-Search, we conduct comprehensive experiments on 4 commercial LLMs with web search, covering 3 test case generation strategies and 5 categories of harmful content. Compared with baselines, CREST-Search achieves a higher risk detection rate of 80.5%. Through detailed analysis, we find that CREST-Search can successfully expose vulnerabilities across all five categories, underscoring the utility of our proposed test case generation strategies. Our ablation study further highlights the contribution of each component of CREST-Search: both the refinement mechanism and the fine-tuned red-teaming model enhance the optimization efficiency and improve the risk detection rate. Our main contributions are as follows:
@@
    We construct a diverse specialized dataset for red-teaming model fine-tuning, named WebSearch-Harm. Compared to the vanilla models, cases generated by the fine-tuned red-teaming models effectively increase the risk detection rate, while reducing the refinement cost and time.
@@
    We propose an innovative red-teaming framework, incorporating three attack strategies, that comprehensively assesses the safety vulnerabilities in the LLMs with web search under the black-box settings.
@@
    By analyzing the vulnerabilities uncovered through our red-teaming framework, we provide suggestions for enhancing the safety of LLMs with web search.
2 Related Work
2.1 LLMs Enhanced by Web Search
Recent advances in large language models (LLMs), such as GPT-4o (OpenAI, 2024a), Gemini (Team, 2025), and LLaMA (Dubey and others, 2024), have significantly improved natural language understanding, demonstrating strong performance across reasoning tasks (Wei et al., 2022; Chen et al., 2021; Kwiatkowski et al., 2019). Beyond NLP, LLMs are increasingly applied in domains such as education (Wang and others, 2023), healthcare (Thirunavukarasu and others, 2023), and finance (Yang et al., 2024). However, LLMs are inherently limited by the fixed cut-off date of their training data. To mitigate this limitation, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) is proposed, which retrieves information from curated external corpora. While this is effective, it often requires costly construction and continuous maintenance of domain-specific knowledge bases.
Alternatively, recent works integrate web search into LLMs, such as GPT-4o-search (OpenAI, 2025) and Gemini-2.5-flash-search (Google, 2025), allowing LLMs to dynamically retrieve information from the open web. Compared to RAG, web-search–augmented LLMs eliminate the need for maintaining specialized corpora. In addition, compared to traditional search engines, LLMs can deeply understand the semantics of user queries and produce concise answers by automatically aggregating and summarizing relevant content across multiple webpages However, this design also introduces new safety and trustworthiness challenges, as malicious webpages may be retrieved and cited as authoritative sources. These risks motivate the need for systematic safety evaluation of LLMs enhanced by web search.
2.2 Adversarial Attacks Against LLMs
@@
Owing to limitations in training data quality and imperfect safety alignment, LLMs remain susceptible to issues such as hallucination, social and cultural bias, and generation of harmful content. Malicious users can craft carefully-designed prompts or poisoning contexts to effectively bypass the safeguard mechanisms of LLMs, thereby misleading them to produce inappropriate content. Numerous attack techniques, such as prompt injection, jailbreak prompting, and adversarial prompt perturbation, have been proposed to exploit these vulnerabilities (Kumar et al., 2023; Yi et al., 2024; Yao et al., 2023). These attacks significantly threaten various real-world applications built upon LLMs.
2.3 Red-teaming LLMs
@@
Building on the above security concerns, red-teaming has emerged as a pivotal strategy for systematically probing the vulnerabilities of LLMs. Red-teaming refers to the deliberate design of adversarial prompts to elicit undesired outputs, thereby stress-testing a model’s safety alignment and uncovering potential risks (Ganguli et al., 2022). This process can provide insightful guidance for further safety refinement (Longpre et al., 2024; Feffer et al., 2024), serving as an essential step before releasing LLM-based products to the public. Researchers have proposed several approaches to red-teaming LLM-based systems (Ganguli et al., 2022; Perez et al., 2022). For instance, Ge et al. (2023) propose an iterative framework, which leverages an adversarial LLM to generate evolving jailbreak prompts. Those prompts, paired with corrected safe responses, are further used to fine-tune the target LLM, achieving scalable and automated safety enhancement for LLMs. An et al. (2025) conduct a systematic analysis towards the safety of RAG for LLMs, demonstrating that even with safe models and safe corpora, RAG can introduce new unsafe behaviors.
@@
Figure 2: Overview of CREST-Search. Different combinations of attack strategies and harmful content categories define distinct attack scenarios. Under each scenario, CREST-Search constructs a unified refinement pipeline consisting of three stages.
3 Preliminary
3.1 Problem Statement